# The Dial-a-Ride Problem with Limited Pickups per Trip

### v2

@@ 5.2 Discussion of the results
In our evaluation of computational performance, we incorporate several performance indicators: time, OBJ, LB, and gap. Time (s) indicates the total time, including the preprocessing time, network time, and CPU time for solver execution. Here, preprocessing time refers to the duration spent on time window tightening and arc elimination, as detailed by Cordeau (2006). Network time refers to the time required for fragment generation and the construction of the fragment-based network. OBJ and LB are the best-found integer solution value and the best lower bound within a 30-minute timeframe, respectively, and gap represents the relative percentage difference between OBJ and LB. When comparing different fragment sets, we also introduce two additional indicators: fragment count and network time. Fragment count is the total number of fragments in the fragment set.
5.2.1 Results for DARP-LPT
Table 5: Computational performance of different formulations for DARP-LPT within 30 minutes
[b] ABF FAF ($S_{FF}$) FFF ($S_{FF}$) Name Time (s) OBJ LB Gap Time (s) OBJ LB Gap Time (s) OBJ LB Gap a3-18 24.3 302.5 302.5 0.00% 0.2 302.5 302.5 0.00% 0.2 302.5 302.5 0.00% a3-30 1800.0 496.7 451.2 9.16% 1800.0 494.8 491.6 0.64% 1800.0 494.8 491.1 0.74% a3-36 1800.0 574.4 541.2 5.78% 1800.0 574.4 565.9 1.49% 1800.0 574.4 565.0 1.63% a4-48 1800.0 712.0 578.3 18.78% 1800.0 698.6 685.4 1.90% 1800.0 697.9 687.4 1.50% a6-60 1800.0 920.2 635.4 30.95% 1800.0 879.2 843.2 4.09% 1800.0 864.9 843.4 2.48% a7-70 1800.0 1284.3 664.6 48.25% 1800.0 1004.0 958.2 4.56% 1800.0 986.6 956.5 3.05% a8-80 1800.0 1359.6 670.4 50.69% 1800.0 1100.7 1017.9 7.53% 1800.0 1071.1 1019.3 4.84% Avg-A 1546.3 23.37% 1542.9 2.89% 1542.9 2.03% b3-36 1800.0 576.5 548.4 4.88% 1800.0 576.5 566.8 1.69% 1800.0 576.5 567.7 1.53% b4-40 1800.0 637.5 612.5 3.92% 1800.0 635.2 630.2 0.79% 1800.0 635.2 631.1 0.64% b5-50 1800.0 796.5 659.3 17.22% 1800.0 798.4 765.7 4.10% 1800.0 789.1 768.3 2.63% b6-60 1800.0 937.9 749.9 20.05% 1800.0 940.6 896.5 4.68% 1800.0 931.7 900.7 3.33% b7-70 1800.0 1162.9 799.9 31.21% 1800.0 965.3 919.4 4.75% 1800.0 950.4 920.7 3.13% Avg-B 1800.0 15.46% 1800.0 3.20% 1800.0 2.25%
@@ 5.2.1 Results for DARP-LPT
    The rows labeled Avg-A and Avg-B represent the average metric values calculated across all instances of type A and type B, respectively.
@@
Table 5 compares the three formulations ABF, FAF, and FFF for DARP-LPT for a runtime limit of 30 minutes, using the basic instances with the parameters $|V|=\lceil n/3\rceil$ and $L=4$. Both FAF and FFF utilize $S_{FF}$ as the input fragment set. The table indicates that, as observed from the time and gap indicators, both FFF and FAF largely outperform ABF, with FFF exhibiting superior performance compared to FAF.
$S_{RF}$ $S_{FF}$ $S_{MF}$ Type F Net Time Gap F Net Time Gap F Net Time Gap Basic, $L=4$ A 475.7 0.4 1542.9 2.17% 567.3 1.4 1542.9 2.03% 452.0 1.7 1543.0 1.92% B 579.6 0.4 1800.0 2.42% 808.0 1.5 1800.0 2.25% 506.0 1.9 1800.0 2.32% Basic, $L=6$ A 475.7 0.4 1107.8 0.86% 937.3 5.8 1034.5 0.68% 454.6 5.8 1057.9 0.65% B 579.6 0.4 1444.6 0.57% 2449.2 10.2 1422.5 0.47% 511.8 11.1 1443.4 0.45% Extended, $L=4$ A 661.0 0.5 1542.9 2.44% 875.3 2.2 1542.9 2.15% 612.7 3.0 1543.0 2.17% B 1914.0 1.7 1800.0 3.67% 3070.0 8.7 1800.0 2.87% 1376.0 8.5 1800.0 2.94% Extended, $L=6$ A 661.0 0.5 1031.5 0.99% 1807.3 15.0 1031.7 0.91% 620.6 14.6 1030.8 0.94% B 1914.0 1.9 1228.9 1.08% 20382.6 204.3 1104.4 1.25% 1495.6 202.8 1191.7 0.96% 1 Type: instance type; F: fragment count; Net: network time, the time required for fragment generation and fragment-based network construction; Time (s) represents the total time, including preprocessing time, network time, and CPU time for solver; basic: basic instances; extended: extended instances.
@@
Table 6: Average computational performance of FFF across all instances using various fragment sets for DARP-LPT within a 30-minute timeframe
@@
Table 6 presents a comparative analysis of three fragment set types $S_{RF}$, $S_{FF}$, and $S_{MF}$. These fragment sets are utilized as input for formulation FFF for solving DARP-LPT. The table covers both basic and extended instances, with $|V|=\lceil n/3\rceil$ when $L=4$ and $|V|=\lceil n/5\rceil$ when $L=6$. Due to the extensive amount of results, we present only the average metrics (fragment count, network time, total time, and gap) for all instances here. Detailed results for basic and extended instances with $L=4$ and $L=6$ are available in Tables 1, 2, 3, and 4 in the Appendix.
@@
We begin by comparing the two traditional fragment sets $S_{FF}$ and $S_{RF}$. The set $S_{FF}$ is better than $S_{RF}$ in most scenarios, particularly excelling in scenarios with a limited number of pickups per trip (e.g., $L=4$), flexible time windows (extended instances), and increased maximum ride time (instance type B), as reflected in the columns Gap. However, for extended type B instances with $L=6$, $S_{RF}$ may outperform $S_{FF}$. This variation can be attributed to fragment count and the quality of lower bound. When $L$ is lower, or when time windows or maximum ride time are tight, the fragment counts for $S_{FF}$ and $S_{RF}$ are similar or at least of the same magnitude. Conversely, when $L$ is higher and both time windows and maximum ride time are more flexible, the fragment count for $S_{FF}$ is substantially higher than that for $S_{RF}$. This is important because employing a large fragment set as input can slow down the computation process. Additionally, the fragments from $S_{FF}$ are longer than those from $S_{RF}$, in a similar way as paths are longer than individual arcs. This similarity helps to explain why $S_{FF}$ generally achieves better lower bounds than $S_{RF}$, because path-based (extended) formulations for routing problems typically yield superior lower bounds compared to arc-based formulations (Desrosiers et al., 2024).
The mixed fragment set $S_{MF}$ demonstrates superior solution quality in scenarios with more flexible time windows and maximum ride time and higher values of $L$ (extended type B instances with $L=6$), as indicated by the gap metrics. Notably, the fragment count of $S_{MF}$ remains relatively low, especially when compared to $S_{FF}$. Since a larger number of fragments can slow down fragment-based formulations, $S_{MF}$ presents itself as an efficient alternative to $S_{FF}$ in this regard.
5.2.2 Results for MDARP-LPT
$S_{RF}$ $S_{FF}$ $S_{MF}$ $|V|$ Type F Net Time Gap F Net Time Gap F Net Time Gap $\lceil n/4\rceil$ A 269.1 8.2 813.9 1.50% 274.4 9.5 806.4 1.19% 162.7 69.7 777.4 0.82% B 1489.2 14.8 1471.1 3.02% 1650.6 20.5 1210.7 1.91% 674.4 96.6 797.5 0.78% $\lceil n/3\rceil$ A 201.9 10.0 636.5 0.56% 206.4 10.9 464.3 0.59% 124.3 52.7 423.5 0.31% B 1195.0 16.7 1107.3 1.80% 1227.6 20.0 872.0 0.55% 518.8 90.9 552.6 0.36% 1 Type: instance type; F: fragment count; Net: network time, the time required for fragment generation and fragment-based network construction; Time (s) represents the total time, including preprocessing time, network time, and CPU time for solver.
Table 7: Average computational performance of FFF across all instances using various fragment sets for MDARP-LPT within a 30-minute timeframe (basic instances with $L=4$)
Table 7 compares three types of fragment sets for MDARP-LPT using FFF under different parameters. Formulation FAF for DARP-LPT can be extended with additional constraints to handle also MDARP-LPT, but its performance significantly lags behind that of FFF, which is why we only present results for FFF and not FAF. The reason for not altering $L$ is that $L=4$ represents a commonly used parameter in carpooling scenarios (the realistic background of MDARP-LPT). Here, we present only the average metrics for all instances. Detailed results can be found in Table 1 and 2 in the Appendix.
Table 7 demonstrates the superiority of $S_{PFF}$ over $S_{RF}$ and $S_{FF}$ in computational performance, particularly for handling large instances. This is evident from the consistently lower time and gap values achieved with $S_{PFF}$ compared to $S_{FF}$. Notably, $S_{PFF}$’s superiority can likely be attributed to its lower fragment count. Our findings also indicate that across most instances, $S_{FF}$ generally outperforms $S_{RF}$ in terms of both time efficiency and solution quality.
@@ 6 Summary and conclusions
To efficiently solve both MDARP-LPT and DARP-LPT, we have applied a unified fragment-based method, where a fragment is a partial path. Our main contributions entail the extensions of a fragment-based formulation, the creation of improved fragment sets, and the insights gained from comparing fragment sets. More specifically, we have modeled (M)DARP-LPT using two fragment-based formulations, FFF and FAF, both of which are extended from the DARP fragment-based formulation of Rist & Forbes (2021), transitioning from fragment-based decision variables to fragment-vehicle-based variables. We find that FFF achieves better results than FAF in terms of solving time and optimality gap, which confirms our theoretical finding that FFF will also yield better lower bounds. We have also introduced improved fragment sets as formulation inputs. The two benchmark fragment sets, the full fragments set $S_{FF}$ and the restricted fragments set $S_{RF}$, are generated using the methods described by Alyasiry et al. (2019) and Rist & Forbes (2021), respectively. Our two novel sets are the path-enumerated full fragment set $S_{PFF}$ and the mixed fragment set $S_{MF}$. The generation of $S_{PFF}$ involves exhaustively enumerating paths, selectively choosing those meeting specific criteria, and then decomposing them into full fragments. This approach draws inspiration from the observation that carpooling and crowdshipping DARP routes typically involve a limited number of customers. We construct the set $S_{MF}$ by decomposing some of the full fragments in $S_{FF}$ into restricted fragments while preserving the remainder. This approach is designed for scenarios when $L$ is higher. All these fragment sets can be employed as input for fragment-based formulations to reach optimality. Our computational results demonstrate that for MDARP-LPT with $L=4$, a realistic carpooling setting, $S_{PFF}$ outperforms $S_{FF}$, and $S_{FF}$ performs better than $S_{RF}$ in terms of computational performance. For DARP-LPT, $S_{FF}$ generally outperforms $S_{RF}$ except when the instances have more flexible time windows, increased maximum ride times, and higher $L$ values. In such cases, $S_{MF}$ can replace $S_{FF}$ to achieve the best performance.
We conclude with two supplementary observations. First, in DARP-LPT, FAF may outperform FFF for a number of small instances, as shown in Table 5. This advantage becomes more prevalent when a fixed number of vehicles is utilized. The underlying reason for this, intuitively, is that FAF, having fewer variables than FFF, solves its linear relaxations more quickly, but produces worse lower bounds. Thus, in small instances with fairly simple branch-and-bound trees, FAF tends to perform better. Conversely, in large instances, FAF’s performance tends to be worse than FFF. Additionally, when the number of vehicles is fixed, the complexity of the problem is reduced, allowing FAF to perform better than FFF in more scenarios. Second, the technique of variable fixing could be a viable option to reduce the number of fragments by removing those whose reduced cost is higher than the gap between a feasible solution’s objective value and the linear relaxation bound (Desaulniers et al., 2020). In the instances that we have tested, however, even if we set the objective value as the optimal value, this method typically removes only a small number of fragments, which hardly affects the final computational results.
Appendix A Arc-based formulation of MDARP-LPT
For MDARP-LPT, the full arc-based formulation (following Cordeau (2006)) is as follows:
Table 1: Sets, parameters, and decision variables of arc-based formulation for MDARP-LPT
Sets Definition
$N_{v}$ Set of locations, $N=P_{v}\cup D_{v}\cup\{o_{v},d_{v}\}$, where $o_{v}$ and $d_{v}$ are respectively
origin and destination of vehicle $v\in V$
$P_{v}$, $D_{v}$ Pickup and delivery locations of vehicle $v\in V$
$A_{v}$ Set of location arcs of vehicle $v\in V$
Parameters Definition
$e_{iv}$, $l_{iv}$ The earliest and latest time of vehicle $v$ at location $i\in N$
Decision variables Definition
$f_{ijv}$ =1, if the arc $(i,j)\in A_{v}$ is traversed by vehicle $v\in V$; =0, otherwise
$t_{iv}$ the departure time of vehicle $v$ on location $i\in N_{v}$
$Q_{iv}$ the load of vehicle $v$ after visiting location $i\in N_{v}$
$\min\sum_{v\in V}\sum_{(i,j)\in A_{v}}C_{ij}f_{ijv}$ (40)
$s.t.$ $\sum_{v\in V}\sum_{j\in N_{v}}f_{ijv}$ $=1$ $i\in P$ (41)
$\sum_{j\in N_{v}}f_{ijv}$ $=\sum_{j\in N_{v}}f_{(i+n)jv}$ $i\in P_{v},v\in V$ (42)
$\sum_{j\in N_{v}}f_{o_{v}jv}$ $=1$ $v\in V$ (43)
$\sum_{j\in N_{v}}f_{jiv}$ $=\sum_{j\in N_{v}}f_{ijv}$ $i\in P_{v}\cup D_{v},v\in V$ (44)
$\sum_{i\in N_{v}}f_{id_{v}v}$ $=1$ $v\in V$ (45)
$t_{iv}+T_{ij}$ $\leq t_{jv}+M_{ijv}(1-f_{ijv})$ $(i,j)\in A_{v},v\in V$ (46)
$e_{iv}\leq t_{iv}$ $\leq l_{iv}$ $i\in N_{v},v\in V$ (47)
$Q_{iv}+q_{i}$ $\leq Q_{jv}+W_{ij}(1-f_{ijv})$ $(i,j)\in A_{v},v\in V$ (48)
$\max(0,q_{i})\leq Q_{iv}$ $\leq\min(Q,Q+q_{i})$ $i\in N_{v},v\in V$ (49)
$T_{i(n+i)}\leq t_{(n+i)v}-t_{iv}$ $\leq R_{i}$ $i\in P_{v},v\in V$ (50)
$t_{(2n+1)v}-t_{0v}$ $\leq R_{0}$ $v\in V$ (51)
$\sum_{(i,j)\in A_{v}}f_{ijv}$ $\leq 2L+1$ $v\in V$ (52)
$f_{ijv}$$\in\{0,1\}$ $(i,j)\in A_{v},v\in V$ (53)
To be concise, we outline the differences between the formulations of MDARP-LPT and DARP-LPT. Constraints (4) to (6) in DARP-LPT are modified to constraints (43) to (45) in MDARP-LPT, accommodating the network flow constraints for distinct vehicles with unique origins and destinations. Constraints (52) can be modified to $\sum_{(i,j)\in A_{v}}\max(q_{i},0)f_{ijv}\\ \leq 2\lambda+1$ if the constraint on the maximum number of pickups per trip is replaced by the constraint on the maximum quantity of items (or passengers) $\lambda$ per trip, where $q_{i}$ represents the customer load quantity. Importantly, the constraints on $\lambda$ remain applicable within our fragment-based approach. The necessary modification involves updating constraint (21) in FFF and constraint (32) in FAF from $\sum_{f\in F}L_{f}x_{fv}\leq L$ to $\sum_{f\in F}\lambda_{f}x_{fv}\leq\lambda$, where $\lambda_{f}$ represents the total quantity of items in fragment $f$.
Appendix B Full computational results for DARP-LPT
All four tables below serve as supplements to Table 6.
$S_{RF}$ $S_{FF}$ $S_{MF}$ Name F Net MIP Time OBJ LB Gap F Net MIP Time OBJ LB Gap F Net MIP Time OBJ LB Gap a3-18 73 0.3 0.0 0.4 302.5 302.5 0.00% 64 0.1 0.0 0.2 302.5 302.5 0.00% 64 0.3 0.2 1.0 302.5 302.5 0.00% a3-30 93 0.1 1799.9 1800.0 494.8 491.0 0.75% 89 0.1 1799.9 1800.0 494.8 491.1 0.74% 89 0.5 1798.8 1800.0 494.8 489.7 1.01% a3-36 121 0.1 1800.5 1800.0 574.4 560.6 2.40% 112 0.2 1797.6 1800.0 574.4 565.0 1.63% 113 0.1 1799.8 1800.0 574.4 566.7 1.35% a4-48 318 0.2 1798.9 1800.0 697.9 686.1 1.69% 312 0.5 1799.4 1800.0 697.9 687.4 1.50% 303 0.5 1798.8 1800.0 697.9 687.9 1.43% a6-60 601 0.4 1793.4 1800.0 869.0 843.2 2.97% 648 1.4 1795.5 1800.0 864.9 843.4 2.48% 579 1.8 1795.7 1800.0 864.9 844.8 2.32% a7-70 1036 0.8 1796.3 1800.0 986.3 958.5 2.82% 1273 2.6 1792.3 1800.0 986.6 956.5 3.05% 932 2.9 1795.1 1800.0 984.9 956.5 2.88% a8-80 1088 0.8 1789.1 1800.0 1064.6 1016.4 4.53% 1473 5.1 1785.7 1800.0 1071.1 1019.3 4.84% 1084 6.2 1782.4 1800.0 1063.4 1016.3 4.43% Avg-A 475.7 0.4 1542.9 2.17% 567.3 1.4 1542.9 2.03% 452.0 1.7 1543.0 1.92% b3-36 169 0.1 1799.9 1800.0 576.5 566.2 1.79% 161 0.1 1799.7 1800.0 576.5 567.7 1.53% 146 0.1 1799.8 1800.0 576.5 566.1 1.81% b4-40 311 0.2 1799.5 1800.0 635.2 630.4 0.76% 203 0.4 1799.1 1800.0 635.2 631.1 0.64% 202 0.4 1799.5 1800.0 635.2 631.8 0.54% b5-50 518 0.4 1798.6 1800.0 789.1 766.2 2.91% 803 1.2 1797.6 1800.0 789.1 768.3 2.63% 496 2.2 1796.2 1800.0 789.5 766.7 2.90% b6-60 940 0.6 1793.3 1800.0 931.4 895.5 3.86% 1218 2.6 1792.3 1800.0 931.7 900.7 3.33% 799 2.0 1794.2 1800.0 931.7 900.4 3.36% b7-70 960 0.7 1796.4 1800.0 948.0 921.6 2.79% 1655 3.2 1789.5 1800.0 950.4 920.7 3.13% 887 4.6 1789.1 1800.0 950.2 921.7 3.00% Avg-B 579.6 0.4 1800.0 2.42% 808.0 1.5 1800.0 2.25% 506.0 1.9 1800.0 2.32% 1 F: fragment count; Net: network time, the time required for fragment generation and fragment-based network construction; MIP: CPU time for solver; Time (s) represents the total time, including preprocessing time, network time, and CPU time for solver. The rows labeled Avg-A and Avg-B represent the average metric values calculated across all instances of type A and type B, respectively.

### v3

@@ 5.2 Discussion of the results
In our evaluation of computational performance, we incorporate several performance indicators: time, OBJ, LB, and gap. Time (s) indicates the total time, including the preprocessing time, network time, and CPU time for solver execution. Here, preprocessing time refers to the duration spent on time window tightening and arc elimination, as detailed by Cordeau [6]. Network time refers to the time required for fragment generation and the construction of the fragment-based network (in FFF and PSFF), or for path generation (in PBF). OBJ and LB are the best-found integer solution value and the best lower bound within a 30-minute timeframe, respectively, and gap represents the relative percentage difference between OBJ and LB. When comparing different fragment sets, we also introduce two additional indicators: fragment count and network time. Fragment count is the total number of fragments in the fragment set.
Table 5: Computational performance of different formulations for DARP-LPT (30-minute limit)
ABF FFF ($S_{FF}$) PSFF ($S_{FF}$) PBF
Name Time OBJ LB Gap Time OBJ LB Gap Time OBJ LB Gap Net CPU Time OBJ LB Gap
a3-18 24.3 302.5 302.5 0.00% 0.2 302.5 302.5 0.00% 0.1 302.5 302.5 0.00% 0.4 0.0 0.5 302.5 302.5 0.00%
a3-30 1800.0 496.7 451.2 9.16% 1800.0 494.8 491.1 0.74% 1.2 494.8 494.8 0.00% 2.5 0.3 3.2 494.8 494.8 0.00%
a3-36 1800.0 574.4 541.2 5.78% 1800.0 574.4 565.0 1.63% 0.4 574.4 574.4 0.00% 6.6 1.3 9.1 574.4 574.4 0.00%
a4-48 1800.0 712.0 578.3 18.78% 1800.0 697.9 687.4 1.50% 0.7 697.9 697.9 0.00% 26.0 3.0 33.2 697.9 697.9 0.00%
a6-60 1800.0 920.2 635.4 30.95% 1800.0 864.9 843.4 2.48% 2.7 864.9 864.9 0.00% 5.2 1.0 7.1 864.9 864.9 0.00%
a7-70 1800.0 1284.3 664.6 48.25% 1800.0 986.6 956.5 3.05% 8.3 984.1 984.0 0.00% 9.9 1.8 13.4 984.1 984.0 0.00%
a8-80 No 1800.0 1071.1 1019.3 4.84% 10.0 1058.7 1058.7 0.00% 25.9 5.6 35.7 1058.7 1058.7 0.00%
a8-96 No 1800.0 1353.6 1302.7 3.76% 13.3 1342.2 1342.2 0.00% No
Avg-A 1504.1 18.82% 1512.9 2.25% 4.59 0.00% 14.6 0.00%
b3-36 1800.0 576.5 548.4 4.88% 1800.0 576.5 567.7 1.53% 0.4 576.5 576.5 0.00% 59.4 12.4 83.0 576.5 576.5 0.00%
b4-40 1800.0 637.5 612.5 3.92% 1800.0 635.2 631.1 0.64% 1.0 635.2 635.2 0.00% 121.0 35.2 174.6 635.2 635.2 0.00%
b5-50 1800.0 796.5 659.3 17.22% 1800.0 789.1 768.3 2.63% 3.1 789.1 789.1 0.00% 239.9 95.8 371.5 789.1 789.1 0.00%
b6-60 1800.0 937.9 749.9 20.05% 1800.0 931.7 900.7 3.33% 4.9 927.1 927.1 0.00% 59.9 25.9 95.8 927.1 927.1 0.00%
b7-70 1800.0 1162.9 799.9 31.21% 1800.0 950.4 920.7 3.13% 7.9 944.7 944.7 0.00% 105.1 33.9 158.0 944.7 944.7 0.00%
b8-80 1800.0 1741.53 889.69 0.49% 1800.0 1104.8 1072.0 2.97% 3.6 1103.2 1103.2 0.00% No
b8-96 No 1800.0 1322.6 1260.9 4.67% 7.4 1294.4 1294.4 0.00% No
Avg-B 1800.0 12.96% 1800.0 2.70% 4.04 0.00% 176.6 0.00%
@@
    Net: network time; CPU: CPU time for solver. The rows labeled Avg-A and Avg-B represent the average metric values calculated across all instances of type A and type B, respectively. “No” indicates that no feasible solution was obtained within the 30-minute time limit.
@@
Table 5 compares the four formulations ABF, FFF, PSFF, and PBF for DARP-LPT for a runtime limit of 30 minutes, using the instances of Cordeau [6] with the parameters $|V|=\lceil n/3\rceil$ and $L=4$. Both FFF and PSFF utilize $S_{FF}$ as the input fragment set. The table indicates that, as observed from the time and gap indicators, FFF and PSFF largely outperform ABF, with PSFF exhibiting superior performance compared to FFF. Even with $L=4$, PBF performs worse than PSFF in terms of the total time. For larger instances (e.g., a8-96 and b8-96), PBF does not have sufficient time to generate all the required paths.
@@
Table 6: Average computational performance of different fragment sets for DARP-LPT and MDARP-LPT (30-minute limit)
$S_{RF}$ $S_{FF}$ $S_{MF}$ $S_{EFF}$ $S_{ERF}$
Problem $L$ Type F CPU T F CPU T F CPU T F CPU T F CPU T
DARP -LPT 4 A 599.4 0.6 1.5 713.0 1.0 4.5 700.3 1.1 3.6 15352.3 1.5 21.5 46594.0 6.6 27.9
B 1231.9 1.9 3.2 1777.7 1.7 9.4 1287.6 1.8 7.9 24920.4 2.7 54.9 103744.7 7.2 55.6
6 A 599.4 9.9 11.2 1153.0 10.8 21.4 707.6 11.1 20.4 29132.4 19.5 69.5 46594.0 207.7 229.9
B 1231.9 9.9 12.0 9905.6 20.5 164.1 1439.0 16.3 170.3 6/7 solved 108323.7 46.7 104.1
MDARP -LPT 4 A 252.5 19.0 35.0 252.5 20.1 39.1 264.9 19.2 36.5
B 2079.6 82.4 124.4 2285.4 47.5 126.9 1708.4 71.4 154.8
DARP -LPT 4 Ried -ler 1301.5 3.4 4.8 2328.4 2.1 6.3 1702.9 2.3 6.2 17471.0 2.6 18.0 60836.1 19.5 46.0
6 1301.5 43.8 45.6 11845.3 49.6 83.2 1782.6 31.0 63.0 73336.5 64.4 172.8 26/30 solved
MDARP -LPT 4 Ried -ler 454.7 17.2 31.6 633.87 14.6 31.9 547.0 14.3 32.0
  • 1
    Type denotes the instance type: A and B refer to type A and type B instances from Ropke & Cordeau [19], respectively; Riedler refers to instances from Riedler & Raidl [16]. F: fragment count; CPU: CPU time for solver; T: Total time (s); “–” denotes that no further testing is performed for this scenario.
@@
Table 6 presents a comparative analysis of the five fragment sets $S_{RF}$, $S_{FF}$, $S_{MF}$, $S_{EFF}$, and $S_{ERF}$ for both DARP-LPT and MDARP-LPT. All these fragment sets are utilized as input for formulation PSFF. The table covers the instances from Cordeau [6] and Riedler & Raidl [16], with varying $L$. Due to the extensive amount of results, we present only the average metrics (fragment count, CPU time, and total time) for all instances here. The total time mainly comprises network time and CPU time, with all other components contributing less than 1%. Since both CPU time and total time are reported, we do not report network time separately. Detailed results for all instances are provided in B.
The table provides several key insights. Our analysis focuses on three main aspects: determining the best fragment set for the current (M)DARP-LPT problem, evaluated by total time; assessing fragment set quality, measured by CPU time alone; and explaining why different fragment sets exhibit varying performance.
Table 6 leads to several noteworthy observations. First, when fed into the formulation PSFF, the fragment set $S_{RF}$ consistently outperforms alternative sets in terms of the total time to reach optimality across all problem instances. Second, when CPU time is considered independently (that is, excluding network time) $S_{FF}$ achieves performance comparable to or even surpassing that of $S_{RF}$ for instances with tight time windows and low values of $L$ (e.g., in DARP-LPT and MDARP-LPT, for instances from Riedler & Raidl [16] with $L=4$). By contrast, in DARP-LPT, for type-B instances with $L=6$, $S_{RF}$ demonstrates a clear advantage over $S_{FF}$, primarily due to the substantially higher fragment count of $S_{FF}$. Third, $S_{MF}$ proves to be an effective alternative of $S_{FF}$ when the fragment count of $S_{FF}$ becomes prohibitively large. For example, for Riedler & Raidl’s instances in DARP-LPT with $L=6$, $S_{MF}$ is the most effective fragment set in terms of CPU time performance. Fourth, both $S_{EFF}$ and $S_{ERF}$ demonstrate inferior performance in terms of the total time to reach optimality, even when considering CPU time alone. This is primarily due to their excessive fragment counts. Analogously to the comparison between $S_{FF}$ and $S_{RF}$, $S_{EFF}$ achieves better performance than $S_{ERF}$ when $L$ is small and the instances have tight time windows and ride times. Given that $S_{ERF}$ and $S_{EFF}$ demonstrate significantly inferior performance compared to other fragment sets in DARP-LPT, we do not further evaluate them for MDARP-LPT.
@@
The performance differences observed across fragment sets can be largely attributed to two factors: fragment count and the quality of the resulting lower bound. The comparison between $S_{FF}$ and $S_{RF}$ is particularly illustrative of this underlying logic. When $L$ is lower, or when time windows or maximum ride time are tight, the fragment counts for $S_{FF}$ and $S_{RF}$ are similar or at least of the same magnitude. Conversely, when $L$ is higher and both time windows and maximum ride time are more flexible, the fragment count for $S_{FF}$ is substantially higher than that for $S_{RF}$. This is important because employing a large fragment set as input can slow down the computation process. Additionally, the fragments from $S_{FF}$ are longer than those from $S_{RF}$, in a similar way as paths are longer than individual arcs. This similarity helps to explain why $S_{FF}$ generally achieves better lower bounds than $S_{RF}$, because path-based formulations for routing problems typically yield superior lower bounds compared to arc-based formulations [8]. Additionally, when transitioning to $S_{MF}$, the set effectively shifts fragments from $S_{FF}$ to ERFs and RFs. As a result, the average fragment length in $S_{MF}$ lies between that of $S_{FF}$ and $S_{RF}$, while the fragment count remains of the same order of magnitude as that of $S_{RF}$. This balance explains why $S_{MF}$ achieves superior CPU time performance in certain scenarios.
Given the relative simplicity of the current problem and the efficiency of the PSFF formulation, the benefits of higher-quality fragments are outweighed by the overhead of fragment enumeration. However, under the slower FFF formulation, the advantage of $S_{FF}$ over $S_{RF}$ becomes more pronounced. For example, in DARP-LPT with $L=4$, using the FFF formulation with a 30-minute time limit (but excluding the instance b8-96, where $S_{RF}$ fails to produce a feasible solution), $S_{FF}$ achieves average optimality gaps of 2.25% (type A) and 2.37% (type B), compared to 2.43% and 2.53% for $S_{RF}$, respectively (see C). Moreover, when the problem becomes more realistic by introducing additional constraints, the fragment enumeration time tends to remain stable (as the number of feasible paths becomes even smaller), whereas the solution time grows exponentially. This underscores the increasing value of employing higher-quality fragment sets as problem difficulty increases. Additionally, even without considering more complex settings but simply increasing the instance size, $S_{FF}$ outperforms $S_{RF}$ in terms of total time; this can be seen, for example, in DARP-LPT with Riedler & Raidl’s largest instance $60N\_5K\_C$ and $L=6$ (where both $S_{MF}$ and $S_{FF}$ even incur higher network construction time). Although the average metrics may not fully capture the importance of fragment set quality, the results on larger instances indicate that more difficult cases favor good-quality fragment sets. One further thing to mention is that some shared mobility studies [26, 23] use EFFs or FFs for DARP or similar problems. While the enumeration time is quite high in our study, it is significantly lower in theirs. This is partly because their instances feature much tighter time windows, and partly because we invest additional time in fragment elimination to ensure the best possible fragment set.
@@ 6 Summary and conclusions
To efficiently solve both MDARP-LPT and DARP-LPT, we have applied a unified fragment-based method, where a fragment is a partial path. Our main contributions entail the extensions of a fragment-based formulation, the creation of improved fragment sets, and the insights gained from comparing fragment sets. More specifically, we model (M)DARP-LPT using two fragment-based formulations, FFF and PSFF, both of which are extended from the DARP fragment-based framework proposed by Rist & Forbes [18]. For FFF the extension resides in the transition from fragment-based to fragment–vehicle–based decision variables, whereas PSFF retains the original variables but augments the network with a pickup-count dimension, resulting in a pickup-space network. We find that PSFF achieves better results than FFF. We also compare several fragment sets used as inputs to the formulations. The full fragment set $S_{FF}$ and the restricted fragment set $S_{RF}$ are generated following the methods proposed by Alyasiry et al. [2] and Rist & Forbes [18], respectively. The extended full fragment set $S_{EFF}$ and the extended restricted fragment set $S_{ERF}$ are constructed based on the approach of Rist & Forbes [17]. In addition, we develop the new mixed fragment set $S_{MF}$ by decomposing full fragments from $S_{FF}$ into (extended) restricted fragments. All these fragment sets can be used as inputs to fragment-based formulations to reach optimality. Our computational results show that, for both DARP-LPT and MDARP-LPT, $S_{RF}$ outperforms $S_{FF}$ in total solution time when used with the PSFF formulation. However, in terms of fragment set quality, measured solely by CPU time, $S_{FF}$ remains superior in instances with tight time windows and a limited number of pickups (which is also close to realistic application scenarios). In DARP-LPT, $S_{MF}$ serves as a high-quality alternative to $S_{FF}$, especially when the size of $S_{FF}$ becomes large.
The incorporation of a time dimension into the formulations FFF or PSFF may be a valuable direction for future work. For example, PSFF can naturally be extended to a pickup-time-space formulation to incorporate both time and pickup dimensions. However, since vehicle indices and pickup count are already used as dimensions, and PSFF performs well under the current setup, we have not pursued this extension in this paper.
Appendix A Additional explanation for fragment-based network
Decomposing feasible DARP routes into fragments yields a significantly smaller number of fragments compared to DARP routes. Conversely, if we combine these fragments, we can obtain all feasible DARP routes, albeit with the possibility of some redundant routes. Fig. A.1 illustrates how feasible DARP routes can be gathered in a transitional network (as an intermediate step towards the fragment-based network). Fig. 1 presents several DARP routes and their corresponding fragments. The network in Fig. A.1 will use these fragments and allow to enumerate all the DARP routes shown in Fig. 1. Identical fragments across different routes in Fig. 1 can be consolidated into a single fragment while maintaining the connections to other fragments. Subsequently, the fragment-based network of Fig. 2 is derived from the network of Fig. A.1. When different fragments share the same start and/or end nodes, then those nodes can also be merged. For instance, two fragments with the same start and end nodes but different route paths $(p1,d1)$ and $(p1,p4,d4,d1)$ become two route paths originating from node $p1$ and terminating at node $d1$. Such consolidations help reduce connections (node arcs). Node arcs with the same start node and end node can also be merged; the three connections from $(p1,d1)$ to $(p2,d2)$, from $(p1,d1)$ to $(p2,p3,d2,d3)$, and from $(p1,p4,d4,d1)$ to $(p2,d2)$ in Fig. A.1, for instance, can be merged into one node arc $(d1,p2)$ in Fig. 2.
Figure A.1: A transitional network for DARP by merging all identical fragments from Fig. 1
Appendix B Full computational results of PSFF for DARP-LPT and MDARP-LPT
The tables below provide supplementary information an the results for PSFF, accompanying Table 6. As nearly all instances achieve optimality across different fragment set inputs, we report only one objective value (OBJ) in the first column and keep four indicators: fragment count, network time, CPU time, and the total time. The test instances comprises: (1) eight Type A and seven Type B instances from Cordeau [2006], and (2) 30 instances from Riedler & Raidl [2018].